Hydrogen exhibits several series of line spectra in different spectral regions. For example the Lyman series ( nf = 1 in Balmer-Rydberg equation) occurs in the ultraviolet region while the Balmer ( nf = 2) series occurs in the visible range and the Paschen ( nf = 3), Brackett ( nf = 4) and Pfund ( nf = 5) series all occur in the infrared range. What is the wavelength (in nm) of a line in the Pfund series where ni = 10?
1
Expert's answer
2020-12-21T03:30:58-0500

ν˜=1/λ=RH(1/n21−1/n22)

Where RH is the Rydberg constant and is equal to 109,737 cm-1.

Pfund series n1=5

n2=10

1/λ=109,737(5-2-10-2) cm-1

      =109,737(0.03)

      =3292.11cm-1

λ = 3.0375655×10-4×107nm

  =3037.57nm